A regular polygon has 39 sides.what is the size of each interior angle?
Answer by Alan3354(69443)

A regular polygon has 39 sides.what is the size of each interior angle?
--------------
The sum = 180*(n-2) = 180*37 = 6660 degs
Each angle = 6660/39 = 2220/13 degs
=~ 170.769 degs
-------------------
Or do it this way.
Each exterior angle = 360/39 = 120/13 degs
Int = 180 - Ext = 2220/13 degs